A project's base case or most likely NPV is $50,000, and assume its probability of occurrence is 60%. Assume the best case scenario NPV is 60% higher than the base case and assume the worst scenario NPV is 30% lower than the base case. Both the best case scenario and the worst case scenario have a 20% probability of occurrence. Find the project's coefficient of variation. Enter your answer rounded to two decimal places. For example, if your answer is 12.345 then enter as 12.35 in the answer box.
Answer #1

Expected NPV=Sum(probability*NPV)=50000*60%+50000*1.6*20%+50000*0.7*20%=53000.00

Standard deviation of NPV=sqrt(Sum(probability*(NPV-Expected NPV)^2))=sqrt(60%*(50000-53000)^2+20%*(50000*1.6-53000)^2+20%*(50000*0.7-53000)^2)=14696.93846

Coefficient of variation=Standard deviation/Expected NPV=14696.93846/53000=0.27730